A ship travels between two ports. The cost of fuel is 100(ax+b/x+10)dollar, where x is the average speed of the ship in knots, and a and b are constants. if the travels at 4 knots, the cost of fuel would be $9000, but at 6 knots, the cost would be $7000.
1) find the values of a and b
2) what is the cost of fuel at an average speed of 5 knots?

Answer by richwmiller(17219)

9000=100((4a+b)/(4+10)),
7000=100((6a+b)/(6+10))
C=100((5a+b)/(5+10))
a = -70, b = 1540, C = 23800/3=7933.33